For what value of k will the graph of kx+6y-2=0 be the same line as the graph of y=2/5x+1/3
Answer by richwmiller(17219)

kx+6y-2=0
y=2/5x+1/3
6y=12x/5+2
6y=-kx+2
k=-12/5